a brick is thrown forward on the ground with a speed of 28 m/s. the coefficient of sliding friction between the brick and the ground is 0.25. compute the time and the distance the brick will travel before coming to rest


1
Expert's answer
2021-05-31T08:39:09-0400

"F_f=ma\to \mu mg=ma\to a=\mu g=0.25\cdot9.8=2.45\ (m/s^2)"


"v=v_0-at\to t=v_0/a=28/2.45=11.43\ (s)"


"s=v_0^2/(2a)=28^2/(2\cdot2.45)=160\ (m)"